Two trains of length 500 metres and 1000 M moving in opposite direction with same speed crosses each other in 10 second,find their speed ?pls explain step by step

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
Relative displacement = 1000 m + 500 m = 1500 m
Relative velocity = v1 + v2 = 2v

Relative velocity = Relative Displacement / Time
2v = 1500 m / 10 s
v = 75 m/s

Speed of each train is 75 m/s